### v2

# Generative Refocusing: Flexible Defocus Control from a Single Image

Abstract
Depth-of-field control is essential in photography, but getting the perfect focus often takes several tries or special equipment. Single-image refocusing is still difficult. It involves recovering sharp content and creating realistic bokeh. Current methods have significant drawbacks. They need all-in-focus inputs, depend on synthetic data from simulators, and have limited control over aperture. We introduce Generative Refocusing, a two-step process that uses DeblurNet to recover all-in-focus images from various inputs and BokehNet for creating controllable bokeh. Our main innovation is semi-supervised training. This method combines synthetic paired data with unpaired real bokeh images, using EXIF metadata to capture real optical characteristics beyond what simulators can provide. Our experiments show we achieve top performance in defocus deblurring, bokeh synthesis, and refocusing benchmarks. Additionally, our Generative Refocusing allows text-guided adjustments and custom aperture shapes. Project page: https://generative-refocusing.github.io/
1 Introduction
Controlling depth of field and focus is essential for artistic photography. It helps guide the viewer’s eye through selective focus and bokeh effects. Getting perfect focus often requires multiple shots or special equipment, which can be tough for everyday photographers. Single-image refocusing removes these obstacles. It allows for focus and bokeh adjustments after the photo is taken. The main challenge is recovering sharp details from blurred areas while creating natural-looking bokeh, while still giving users fine control.
Existing research addresses different aspects of this problem. Defocus deblurring methods [1, 8, 36, 56, 67, 73, 84] aim to recover sharp images. Diffusion approaches [41, 42] show promise for addressing spatially variant blur. Bokeh synthesis methods [9, 30, 48, 60] focus on rendering realistic depth-of-field. Recent efforts include physics-based [38, 37, 50], neural [48, 49], and diffusion-based [19, 82, 71] approaches. However, most methods handle either deblurring or bokeh synthesis individually. End-to-end refocusing often requires specialized capture setups [40, 46, 2].
Despite these advances, three main limitations remain (Tab. 1). First, most methods rely on all-in-focus inputs and accurate depth maps. This restricts their use to images that already have defocus blur. Second, while synthetic data [48, 89] can be used for training, it is limited by the realism of the simulators. This approach often fails to capture detailed visual appearances. On the other hand, obtaining real paired datasets [60] is very expensive. Third, current methods usually only support aperture size, not its shape. This restriction limits the possibilities for creative bokeh effects.
Table 1: Comparison across methods.  = supported;  = not $\LARGE\triangle = supported$ with caveat. In the synthetic/real training data columns, “paired” indicates the use of paired supervision, while “unpaired” indicates real unpaired data (no real paired supervision). Bokehlicious typically assumes an all-in-focus input and requires fine-tuning for different input types. For BokehMe, only the classical renderer variant supports aperture-shape control. UW = ultra-wide, W = wide.
Method Input Training data Supported user control
Setup AIF required? Synthetic Real Aperture size Aperture shape Focus point Defocus map
[…]
DC2 [2] Dual-view (UW+W) paired
Bokehlicious [60] Single-view paired
BokehDiff [89] Single-view paired
DiffCamera [71] Single-view paired
Ours Single-view paired unpaired
To tackle these issues, we present Generative Refocusing, a flexible single-image refocusing system built on a two-stage design. Our DeblurNet module produces a sharp, in-focus image from a variety of inputs. It uses a diffusion model guided by initial deblurring predictions. Our BokehNet creates fully customizable bokeh, considering user-defined focus planes, bokeh intensity, and aperture shapes (LABEL:fig:teaser). A significant advancement in our method is our training scheme, which makes use of both 1) paired synthetic data to maintain geometric consistency and 2) unpaired real-world bokeh images with camera metadata that help us learn real lens characteristics not captured by simulators. Our approach achieves top performance in defocus deblurring, bokeh generation, and refocusing, while maintaining the natural consistency of scenes.
Our main contributions are:
  • A flexible refocusing pipeline accepts images in any focus state. It gives users control over the focus plane, bokeh intensity, and aperture shape through two-stage decomposition.
  • This semi-supervised framework connects synthetic and unpaired real bokeh images. It enables genuine optical learning that goes beyond what a simulator can offer. This is achieved through parameter estimation from EXIF metadata.
  • The system performs well across defocus deblurring, bokeh synthesis, and refocusing. This has been validated on existing benchmarks (DPDD [1], RealDOF [36]) and our new light-field datasets (LF-Bokeh, LF-Refocus). It also has applications in text-guided deblurring and creating aperture shapes.
2 Related Work
Figure 2: Pipeline Overview. Our method decomposes single-image refocusing into two stages: (a) Defocus Deblurring and (b) Shallow Depth-of-Field Synthesis. Given a blurry input image $I_{\text{in}}$, we optionally apply a pre-deblurring method [56] to obtain a conservative estimate $I_{\text{syn}}$, then feed both $I_{\text{in}}$ and $I_{\text{syn}}$ into DeblurNet to recover a high-quality all-in-focus image $I_{\text{aif}}$. The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ convert images to latent representations for processing by the DiT backbone. In the second stage, BokehNet takes the all-in-focus image $I_{\text{aif}}$, the defocus map $D_{\text{def}}$, and optionally a specific aperture shape as inputs to synthesize the refocused output. The defocus map $D_{\text{def}}$ is computed based on the estimated depth map $D$ [7], along with both the user-specified focus plane $S_{1}$ and bokeh level $K$.
Diffusion Models for Image Restoration.
Diffusion models [24, 65, 81, 68] advance image restoration via generative priors, extending to zero-shot restoration and artifact removal. Methods evolved from pixel-space [57] to efficient latent-space approaches [54, 76], enabling faster inference. ResShift [83] cuts steps from 1000 to 4–15 via residual shifting. Deblurring progressed from merging diffusion priors with regression tasks, such as matting [74] or hierarchical restoration [13], to learning spatially-varying kernels in latent space [33]. Cascaded pipelines [25] refine at multiple resolutions, while principled frameworks [45, 44] treat degradations as stochastic processes. Training-free methods [91, 32] enable plug-and-play restoration. Unlike these general methods, we target spatially-varying defocus blur with a two-stage pipeline explicitly separating deblurring from bokeh synthesis for controllable refocusing.
Defocus Deblurring.
Defocus deblurring recovers sharp images from spatially varying blur caused by limited depth of field. Early deconvolution [4, 92] introduced artifacts at depth transitions. Dual-pixel sensors [1], quad-pixel data [8], and disparity-aware techniques [36, 79, 70] improved blur–depth modeling and the recovery of all-in-focus images from focal stacks. Architectures shifted from CNNs [10] to transformers [67, 73], with gains from implicit representations [51, 17] and multi-scale attention [87]. Recent work includes vision-language fusion [77] for semantic estimation and a shift from supervised [52] to diffusion-based learning [42, 41], including unpaired data. While these excel as standalone solutions, we position deblurring as the first stage of refocusing, using FLUX’s [61] generative prior to enable controllable bokeh synthesis from any input.
Bokeh Rendering.
Bokeh rendering progressed from physically based scattering [50, 38, 37] and differentiable rendering [63] to neural methods combining ray tracing with learned enhancements [48, 49]. Learning-based approaches evolved from fixed apertures [30] to variable f-stops [60, 9], extending to video [80] and 3D scenes [62, 72]. Diffusion models enabled bokeh control [19, 82], mainly for text-to-image synthesis. These methods require all-in-focus inputs and accurate depth. We remove this constraint via flexible input handling, which accepts defocused or all-in-focus images, and semi-supervised learning on unpaired real-world bokeh to capture authentic lens characteristics beyond simulators.
Single-Image Refocusing.
Post-capture refocusing manipulates focus after capture. Light-field cameras [46, 40] enable optical refocusing but need specialized capture. Computational approaches evolved from deconvolution [4, 86] producing artifacts, to GAN-based paradigms [58], to diffusion methods [71]. 3D representations [62, 72, 35, 43] achieve real-time controllable DoF but need multi-view capture or calibration. We combine three capabilities: (1) semi-supervised learning mixing synthetic and unpaired real data for authentic optics, (2) single-image flexibility accepting any input without preprocessing, and (3) comprehensive control over focus, intensity, and aperture shape.
Camera-Conditioned Diffusion.
Conditioning diffusion models on camera parameters progressed from extrinsic control for view synthesis [26, 14] and video generation [22, 75], to intrinsic control of focal length, aperture, and exposure [18, 12]. Direct conditioning often yields inconsistent content. Recent work uses temporal modeling [82], explicit calibration [6, 16], or transformer-based pose handling [3, 59]. These focus on generation; we operate in editing, manipulating existing images with scene consistency. By decoupling into deblurring and bokeh stages, we achieve efficient single-image manipulation without multi-frame overhead.
Semi-Supervised Learning.
When paired data is scarce, semi-supervised learning provides alternatives. The field evolved from learning without clean targets [39, 34, 27] to pseudo-labeling [11, 23, 69]. Foundation models enabled the generation of high-quality pseudo-labels [21, 85]. Unpaired frameworks [31, 90, 64] learn enhancement without paired supervision. Two-stage approaches [29, 78] bridge fidelity and perceptual quality via different supervision types. We adapt these to bokeh rendering: perfect paired refocusing data is nearly impossible to obtain. We train on synthetic pairs for structure, then unpaired real bokeh for realism. Deblur restores to all-in-focus, enabling bokeh to learn authentic lens characteristics exceeding simulators.
3 Method
Single-image refocusing couples two operations: recovering sharp content in out-of-focus regions and applying controllable bokeh to originally sharp areas. Blur magnitude is dictated by a defocus map parameterized by a user-specified focus point and the scene depth. When the input is blurry, monocular depth estimation is brittle, undermining precise control over the defocus map and often yielding misfocus or artifacts. We propose Generative Refocusing, which decomposes the task into defocus deblurring and bokeh synthesis, and comprises two models, DeblurNet and BokehNet, that jointly enable precise and controllable refocusing (see Fig. 2 for an overview of the pipeline); Such a flexible pipeline allows us to exert greater control over input conditions (see Tab. 1).
Figure 3: Training data generation. Each training sample consists of the following five components: (i) a bokeh image, (ii) an all-in-focus (AIF) image, (iii) a depth map $D$, (iv) a bokeh level $K$, and (v) a focus plane $S_{1}$. We construct these samples via three routes: (a) Synthetic paired data. Given real AIF images and depth maps $D$, we compute a defocus map $D_{\text{def}}$ parameterized by a specified bokeh level $K$ and focus plane $S_{1}$, and feed it into a bokeh renderer [48] to synthesize corresponding bokeh images. (b) Real unpaired data. Given real bokeh images, DeblurNet recovers an AIF image. We then estimate depth and extract a foreground mask [88] to define the focus plane $S_{1}$. The bokeh level $K$ is computed from the EXIF metadata following the formulation in [19]. (c) Real paired data without EXIF. For real pairs lacking EXIF, we obtain $S_{1}$ as in (b), and follow Eq. (2) to estimate the bokeh level $K$.
3.1 Stage 1: Defocus Deblurring
DeblurNet. We condition the deblurring network on a pair $(I_{\text{in}},I_{\text{pd}})$, where $I_{\text{pd}}$ is a conservative pre-deblur generated by an off-the-shelf defocus deblurring model. In our implementation, we use a non-generative defocus model  [56], whose outputs are typically content-faithful yet over-smoothed. Our model, guided by a learned diffusion prior, then injects high-fidelity details without sacrificing geometric correctness.
Artifact-robust dual conditioning.
To prevent artifacts in $I_{\text{pd}}$ from biasing the reconstruction and to ensure that fine details are recovered from $I_{\text{in}}$, we introduce two complementary mechanisms (see Fig. 2 (a)):
  • Positional disentanglement. We encode the two inputs with disjoint 2D positional grids: $I_{\mathrm{in}}$ is anchored at $(0,0)$, while $I_{\mathrm{pd}}$ is shifted to start at $(0,N)$, where $N$ denotes the side length of the token grid after downsampling $I_{\mathrm{in}}$. This shift lets the network distinguish the two sources, using $I_{\mathrm{pd}}$ as a low-frequency anchor while extracting high-frequency evidence from $I_{\mathrm{in}}$.
  • Pre-deblur dropout. During training, we stochastically replace $I_{\mathrm{pd}}$ with a zero tensor (at a fixed probability), forcing the network to remain effective when the auxiliary input is unreliable or absent. This regularizes the fusion, shields against method-specific artifacts, and enables the model to operate in a single-input deblurring mode at test time.
3.2 Stage 2: Shallow Depth-of-Field Synthesis
Synthetic-pair learning. BokehNet expects an AIF image, a target bokeh image, and the corresponding defocus map (see Fig. 2 (b)). In practice, perfectly paired data and accurate defocus maps are hard to acquire. We therefore pretrain with synthetic pairs: starting from real all-in-focus images and their estimated depth, we randomly sample a focus plane $S_{1}$ and a target bokeh level $K$, and use a simulator [48] to render bokeh consistent with the induced defocus map (see Fig. 3 (a)). This stage quickly trains the network to modulate the circle of confusion based on the defocus map, but it remains constrained by renderer bias and may introduce non-realistic artifacts.
Real unpaired bokeh learning. To capture real optics, we further leverage unpaired real bokeh images with EXIF metadata (see Fig. 3 (b)) and estimate an approximate bokeh level K\!K following [19]:
$K\;\approx\;\frac{f^{2}\,S_{1}}{\,2\,F\,(S_{1}-f)\,}\times\textit{pixel\_ratio},$ (1)
where $f$ is the focal length and $F$ denotes the aperture $f$-number; and pixel_ratio compensates for sensor size and image resolution differences. The AIF input is provided by DeblurNet. Although $S_{1}$ can sometimes be read from metadata, both estimated and true depths may deviate in real scenes. To obtain more reliable supervision, we estimate in-focus masks using BiRefNet [88], and have human experts filter erroneous cases and complete partial masks. These masks refine the focus plane and yield a consistent defocus map, enabling BokehNet to learn from real bokeh images.
Real paired bokeh images without metadata.
Previous bokeh-synthesis datasets [60] [30] provide real pairs but omit EXIF metadata or provide insufficient fields to estimate a reliable $K$ (see Fig. 3 (c)). We therefore adopt a simulator-in-the-loop calibration of the bokeh level: given an AIF image $I_{\mathrm{aif}}$, an estimated depth $D$, and a focus annotation obtained as in the unpaired setting (BiRefNet-based in-focus masks with expert filtering), we sweep $K$ and choose the value whose rendered result best matches the real bokeh target $I_{\mathrm{real}}$:
$K^{\star}=\operatorname*{argmax}_{\begin{subarray}{c}K\in\\ (K_{\min},\,K_{\max})\end{subarray}}\operatorname{SSIM}\!\big(B(I_{\mathrm{aif}},D;\,S_{1},K),\,I_{\mathrm{real}}\big),$ (2)
where B denotes our physically guided renderer and $S_{1}$ is the focus-plane proxy. The selected $K^{\star}$ is then used as the pseudo bokeh level label for training . Focus-plane annotation follows the same protocol as in the unpaired case.
Discussion.
Although the resulting defocus maps and the AIF proxy are not perfectly accurate, this supervision regime mirrors the practical refocusing scenario. Aggregating numbers of real bokeh images substantially improves realism and robustness over synthetic-only training.
3.3 Bokeh-Shape Aware Synthesis
State-of-the-art learning-based bokeh synthesis methods deliver compelling results but do not expose aperture-shape control. Physics-based renderers, in principle, support arbitrary shapes, yet public implementations typically omit this functionality. We therefore explore explicit shape-aware control using only a bokeh shape image within a trainable framework.
Table 2: Defocus deblurring benchmark on RealDOF [36] and DPDD [1] datasets. We evaluate the task of recovering all-in-focus images from defocused-blur inputs. Our method, GenRefocus, uses the DeblurNet module for this task. We report both reference-based metrics (LPIPS, FID) and no-reference perceptual quality metrics (CLIP-IQA, MANIQA, MUSIQ). Best, second best, and third best results are highlighted. GenRefocus outperforms all competing approaches across both benchmarks and all metrics, demonstrating the effectiveness of our diffusion-based deblurring with pre-deblur conditioning.
RealDOF [36] DPDD [1]
Method LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
Input 0.5241 92.93 0.3562 0.2213 28.7087 0.3485 88.85 0.4337 0.3325 45.5376
GT 0.4233 0.2675 42.5777 0.4563 0.3294 47.8102
AIFNet [55] 0.4132 72.38 0.3814 0.2536 35.7121 0.3090 79.84 0.4359 0.3371 46.9558
IFANet [36] 0.2994 33.21 0.4088 0.2609 39.5949 0.2301 42.60 0.4472 0.3276 47.2784
DRBNet [56] 0.2550 29.89 0.3889 0.2609 38.9014 0.1819 39.37 0.4142 0.3254 46.7229
Restormer [84] 0.2863 29.80 0.4062 0.2642 40.8355 0.1762 33.45 0.4332 0.3307 47.9872
INIKNet [51] 0.2851 32.18 0.3984 0.2580 38.9959 0.1860 36.63 0.4290 0.3265 47.1961
GenRefocus (Ours) 0.2356 24.73 0.4575 0.2886 43.7042 0.1598 33.08 0.4619 0.3411 49.3019
Figure 4: Qualitative comparison on defocus deblurring. Visual results on (a) RealDOF [36] and (b) DPDD [1] datasets. Blue boxes on the left indicate cropped regions shown in detail. Our DeblurNet faithfully recovers fine text details (“NEW YORK”) in (a) where competing methods produce blurry or distorted results. In the challenging example (b) with severe defocus blur, other methods either fail to recover structure (AIFNet [55], DRBNet [56], INIKNet [51]) or introduce artifacts in the background (IFANet [36], Restormer [84]), while our method reconstructs geometrically consistent, visually compelling content by leveraging diffusion priors guided by the pre-deblurred input.
Simulator and Data.
Real photographs exhibiting diverse bokeh shapes are rare, and paired AIF–bokeh examples are even scarcer. We thus rely on simulation. A key observation is that when an all-in-focus (AIF) image lacks point-light stimuli, the simulated bokeh carries weak shape cues, making the model reluctant to learn shape-conditioned responses. To address this, we synthesize a PointLight-1K AIF set designed to reveal aperture shape. We also extend the classical BokehMe renderer to scatter through a given shape: given an AIF image $I_{\mathrm{aif}}$, depth $D$, focus-plane proxy $S_{1}$, bokeh level $K$, and a shape kernel $s$ (binary/raster PSF), the simulator renders
$I_{\mathrm{syn}}\;=\;\mathcal{R}\!\big(I_{\mathrm{aif}},D;\,S_{1},K,s\big),$ (3)
yielding paired supervision for shape-aware training.
Shape-conditioned fine-tuning.
Starting from our base model trained without shape control, we freeze all original LoRA weights and introduce another conditioning branch that ingests a shape exemplar $s$. Only this branch is fine-tuned while the backbone remains fixed, ensuring that shape edits do not regress the learned bokeh synthesis.
4 Experiments
4.1 Setup
Backbone and training details.
Our base generative backbone is FLUX-1-dev [61], fine-tuned via LoRA [28] with a multi-condition scheme following [66]. DeblurNet employs LoRA rank r=128r{=}128, while BokehNet uses r=64r{=}64.
Both models are trained with per-GPU batch size $1$ and gradient accumulation over $8$ steps on $4{\times}$RTX A6000 GPUs. DeblurNet is trained for $60\mathrm{K}$ steps. BokehNet is trained in two stages: synthetic pair training for $40\mathrm{K}$ steps and Real unpair for another $40\mathrm{K}$ steps.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input GT BokehMe [48] Bokehlicious [60] BokehDiff [89] Ours
Figure 5: Qualitative comparison on bokeh synthesis benchmark. Results on LF-Bokeh with zoomed patches (blue and orange boxes) highlighting detail quality. Our BokehNet synthesizes bokeh effects that better match ground truth with realistic blur gradients and natural occlusion handling. Baselines show various artifacts: BokehMe [48] exhibits simulator bias, Bokehlicious [60] over-smooths details, and BokehDiff [89] produces inconsistent defocus. Our semi-supervised training on real bokeh images enables the capture of authentic lens characteristics.
Datasets.
We summarize the training data for the two modules below.
DeblurNet. We train on DPDD [1] together with a curated subset of RealBokeh_3MP [60], totaling approximately 4.5K image pairs.
BokehNet (synthetic pairs). We use AIF images from [82] and a subset of EBB! [30], and procedurally render defocus under varying aperture sizes/shapes and focus points to generate about 70K synthetic pairs.
BokehNet (real supervision). We incorporate (i) an unpaired real-bokeh set collected from Flickr by Fortes et al. [19], and (ii) real AIF–bokeh pairs with incomplete metadata from RealBokeh_3MP [60] and LFDOF [55], contributing approximately 27K real examples in total. All datasets are used strictly for research and are split so that no scene overlaps across training and evaluation.
Table 3: Bokeh synthesis benchmark. Evaluation on LF-Bokeh (200 light-field images with multiple focus planes). Following [89], we use per-image binary search for optimal $K$. Bokehlicious supports only aperture size control. Our semi-supervised training on real bokeh data outperforms all baselines.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
BokehMe [48] 0.1228 0.0744 0.9511
Bokehlicious [60] 0.1799 0.1062 0.9304
BokehDiff [89] 0.1708 0.0933 0.9192
GenRefocus (Ours) 0.1047 0.0611 0.9570
Benchmarks.
We evaluate three tasks: (i) defocus deblurring on DPDD and RealDOF; (ii) bokeh synthesis on our new LF-Bokeh benchmark, comprising $200$ images with multiple focus planes synthesized from [53] and [15] light-field captures; and (iii) refocusing on LF-Refocus, built by pairing same-scene images within LF-Bokeh to form $400$ source–target pairs.
Inference-time resolution and aspect ratio.
We adopt a tiling strategy inspired by  [5] at evaluation time: we split the input image into overlapping patches and perform tile-wise denoising before blending them back together. As a result, all benchmarks are conducted at each dataset’s original aspect ratio and resolution, without additional resizing.
4.2 Comparison
Defocus deblurring.
As stated on Tab. 2, DeblurNet consistently outperforms all baselines across reference and no-reference metrics. On certain NR-IQA measures, it even surpasses the ground-truth. Qualitatively, Fig. 4 (a) shows faithful recovery of the “NEW YORK” text with well-preserved edges. The scene in Fig. 4 (b) is particularly challenging: competing methods recover only a blurry dark stripe and further distort the background, whereas DeblurNet reconstructs structure-consistent content and delivers visually compelling results without introducing artifacts or geometry drift.
Table 4: Quantitative comparison on refocusing benchmark. Results on our LF-Refocus dataset (400 pairs with different focus planes). We compare our DeblurNet+BokehNet pipeline against combinations of existing all-in-focus estimation methods [56, 84] and bokeh synthesis methods [48, 89]. Optimal bokeh level $K$ is determined via per-image binary search following [89]. Our approach achieves superior performance across all metrics, demonstrating the benefit of joint optimization and semi-supervised training.
AiF Estimation Bokeh Synthesis LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ NIQE $\downarrow$ CLIP-IQA $\uparrow$
DRBNet [56] BokehMe [48] 0.1471 0.0937 0.9335 43.3087 6.5783 0.4460
BokehDiff [89] 0.1848 0.1030 0.9310 48.0766 5.8195 0.4997
Restormer [84] BokehMe [48] 0.1554 0.0963 0.9303 43.3188 7.1249 0.4782
BokehDiff [89] 0.1860 0.1016 0.9320 48.0326 6.3216 0.5352
DeblurNet (Ours) BokehNet (Ours) 0.1458 0.0850 0.9451 52.6954 5.3343 0.5763
Bokeh synthesis.
Following the BokehDiff protocol [89], we conduct a per-image binary search over the bokeh level $K$ and select the value that maximizes SSIM with the target. All other control signals are held fixed, ensuring a fair comparison in which each method operates at its best attainable setting. As summarized in Tab. 3, our approach attains the highest scores on both fidelity and perceptual metrics, outperforming all baselines. The qualitative results in Fig. 5 show that our method synthesizes bokeh that more closely matches that of real photographs. In the zoomed-in patches, our outputs more faithfully reproduce the ground-truth bokeh than competing baselines. Other methods tend to naively translate the predicted defocus map into blur strength, so any error in the depth scale directly leads to over- or under-blurring. In contrast, BokehNet, trained with our semi-supervised strategy, is more robust to such depth-scale inaccuracies and thus produces a more appropriate amount of blur that better matches the ground truth.
Refocusing.
We build four two-stage baselines by pairing strong AIF estimators (DRBNet [56], Restormer [84]) with bokeh synthesizers (BokehMe [48], BokehDiff [89]), yielding the $2\times 2$ combinations. For each method, we also follow the protocol of BokehDiff [89] to select the best bokeh level $K$ . As shown in Tab.4, our approach achieves the best results on both fidelity and perceptual metrics, outperforming all baseline pairings. Qualitatively, Fig. 6 shows that competing methods tend to produce spatially ambiguous blur and fail to place the focal plane correctly, resulting in uniformly soft content and inaccurate foreground/background separation. In contrast, our approach restores fine details at the intended focus while synthesizing physically consistent bokeh elsewhere.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input GT DRBNet [56] DRBNet [56] Restormer [84] Restormer [84] GenRefocus
+ BokehMe [48] + BokehDiff [89] + BokehMe [48] + BokehDiff [89] (Ours)
Figure 6: Qualitative comparison on refocusing benchmark. Results on LF-Refocus comparing our end-to-end trained pipeline against baseline combinations of existing deblurring [56, 84] and bokeh synthesis [48, 89] methods. Zoomed patches (blue and orange boxes) highlight focal and defocused regions. Baselines struggle with accurate focus-plane placement and produce inconsistent or overly smooth blur. Our method achieves sharp refocusing at the target plane while synthesizing photorealistic bokeh, closely matching ground truth across diverse scenes.
Table 5: Ablation on pipeline design. Comparing one-stage (direct refocusing) and two-stage (deblur then re-bokeh) on LF-Refocus. Both use the same backbone and synthetic data. Two-stage enables better depth control and independent semi-supervised learning per stage, yielding substantial improvements.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
One stage 0.1723 0.0976 0.9345
Two stage (ours) 0.1458 0.0850 0.9461
Table 6: Ablation on BokehNet training. Comparing synthetic-only vs. adding real unpaired bokeh images on LF-Bokeh. Second stage uses real bokeh with estimated parameters (Sec. 3.2), initialized from synthetic weights. Real data enables learning authentic optical characteristics beyond simulators.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
Synthetic pair 0.1289 0.0738 0.9461
Synthetic pair+ Real unpaired 0.1047 0.0611 0.9570
4.3 Ablation Study
One stage and two stage refocusing.
We also evaluate a variant that takes an input image and a provided defocus map and directly produces a refocused output in a single pass. Concretely, we instantiate a one-stage model by reusing the BokehNet backbone and training it on the same synthetic pairs, then assess it on the refocusing benchmark. As reported in Tab. 5, the one-stage variant lags the proposed two-stage pipeline by a substantial margin. We attribute this gap to two factors. First, control fidelity: the defocus map relies on depth estimated from a blurry input; without an all-in-focus view, depth quality degrades, leading to misaligned or biased blur fields and weakening refocusing accuracy. Second, data leverage: the one-stage formulation cannot separate the problem into defocus removal and bokeh synthesis, preventing us from injecting real-data supervision tailored to each subtask.
Key idea and BokehNet training ablation.
Our core idea is to use data generation to harvest real-world supervision for training BokehNet, leveraging real unpaired images. To quantify its impact, we evaluate two training settings without introducing additional architectural changes: (i) synthetic pair: BokehNet is trained purely on synthetic paired data; and (ii) synthetic pair + real unpaired: starting from (i), we further train with real unpaired images processed by our pipeline to produce usable supervision. On the bokeh-synthesis benchmark, the synthetic pair setting performs on par with a classical simulator baseline, while the synthetic pair + real unpaired setting delivers substantial improvements across both fidelity and perceptual metrics (Tab. 6).
Application.
As illustrated in Fig. 7, after we fine-tune BokehNet on our point-light 1K dataset, the model can synthesize bokeh with diverse, user-specified aperture shapes. Moreover, even though DeblurNet is trained without any text prompts, it can still leverage textual instructions at inference time to refine the all-in-focus output, as shown in Fig. 8. This highlights a key advantage of generative defocus-deblurring models: they naturally support post-hoc, prompt-based correction and editing of the reconstructed content.
Refer to caption Refer to caption Refer to caption Refer to caption
Input Triangle Heart Star
Figure 7: Controllable aperture shape synthesis. Given an input image, our BokehNet can synthesize bokeh effects with custom aperture shapes (triangle, heart, star) by conditioning on shape exemplars. Background point lights exhibit the specified aperture geometry while maintaining scene consistency.
Figure 8: Text-guided deblurring. DeblurNet uses optional text prompts to guide reconstruction of severely blurred text. With prompts, we correct hallucinations (red: “DESION” $\rightarrow$ “DESIGN”) and recover accurate text (green: “yeurs” $\rightarrow$ “years”), matching GT. Without guidance, it hallucinates incorrect content.
5 Conclusion
We present Generative Refocusing, a two-stage diffusion framework turning single images into controllable virtual cameras. By decoupling into DeblurNet (deblurring) and BokehNet (bokeh synthesis), our approach handles arbitrary inputs with intuitive controls over focus plane, bokeh intensity, and aperture shape. Through semi-supervised training on synthetic and real unpaired bokeh data, GenRefocus learns authentic optical behavior beyond simulators. Experiments on DPDD, RealDOF, LF-Bokeh, and LF-Refocus show consistent improvements, enabling text-guided refinement and aperture-shape editing.
Limitations and future work.
GenRefocus relies on monocular depth estimation; severe failures result in mislocalized focal planes. Complex aperture shapes require simulator-driven training data. Future work should generalize to richer aperture vocabularies, including user-drawn designs.
Appendix A Appendix Overview
In this appendix, we provide additional details on the training datasets used in our framework, the construction pipeline of our PointLight-1K dataset, a comparison with the VLM model, and visual comparisons from the BokehNet training ablation study. We also present failure cases and additional in-the-wild results on the accompanying HTML page.
Figure 9: Comparison with VLM model. Qualitative refocusing results comparing our Generative Refocusing framework with Gemini 3 Nano Banana Pro [20] given prompt ”focus on the man on the right”.
Appendix B Visual Comparison for BokehNet Training Ablation
In the main paper, we compare different training strategies for BokehNet. As illustrated in Fig. 10, under a foreground-bokeh setting with comparable bokeh strength in both rows, the model in (a), which is trained only on synthetic pairs, inherits artifacts from the simulator [48] and fails to correctly blur fine structures such as hair along object boundaries. In contrast, the model in (b), which is additionally trained with unpaired real bokeh data, produces smoother and more realistic blur around these edges.
Appendix C Training Datasets
We summarize the datasets used to train DeblurNet and BokehNet.
C.1 DeblurNet
For training DeblurNet, we use the DPDD [1] dataset and adopt all images from its official training split. This dataset provides paired defocused and all-in-focus images, which we use to supervise the deblurring stage.
We further leverage the RealBokeh_3MP [60] dataset, which contains images captured with relatively small apertures and thus does not strictly provide all-in-focus views. To obtain images that are sufficiently sharp for our purposes, we compute the Laplacian variance of each image as a simple focus measure and rank the dataset accordingly. We then retain the top $3000$ images with the highest Laplacian variance as additional training data for DeblurNet.
C.2 BokehNet
For training BokehNet, we require diverse all-in-focus inputs. We therefore draw all-in-focus images from the [82] and [30] collections. As in the DeblurNet stage, we compute the Laplacian variance for each candidate image and filter out blurry examples. After this filtering, we obtain a pool of approximately 1.71.7k all-in-focus images for simulating bokeh images. The handling of the real bokeh datasets BokehDiffusion [19] and RealBokeh_3MP [60] for bokeh synthesis is already described in the main paper.
Appendix D PointLight-1K Collection Pipeline
We construct the PointLight-1K dataset to better model scenes with strong point-light bokeh. The pipeline consists of the following steps:
  1. 1.
    Keyword mining. We first mine keywords from photographs on Flickr [^1] https://www.flickr.com/ that are tagged with terms such as “night” or “bokeh”. These keywords capture typical compositions and scene elements with prominent point lights.
  2. 2.
    Prompt expansion. The collected keywords are then expanded into diverse, natural-language prompts using GPT-4o [47]. This step enriches the textual descriptions while preserving point-light characteristics.
  3. 3.
    Image generation with fine-tuned FLUX. We use a fine-tuned version of FLUX.1-Dev [61] (FLUX.1-dev-LoRA-AntiBlur LoRA) to generate images from the prompts. The fine-tuning is designed to encourage sharper scenes.
  4. 4.
    Deblurring with DeblurNet. Despite fine-tuning, we observe that images involving point lights may still exhibit mild residual blur. To further reduce this blur, we pass all generated images through our DeblurNet module, optimizing each image toward an all-in-focus appearance.
  5. 5.
    Final dataset. After this process, we obtain a curated set of 11k night-time images with prominent point lights and minimal residual blur. We refer to this dataset as PointLight-1K and use it as a specialized training set to pretrain and refine our learning strategy for controllable bokeh shape.
Appendix E Comparison with VLM model
We also compare our method with the recently released vision–language model Gemini-3 Nano Banana Pro [20], which explicitly supports text-driven refocusing. For this model, we use the prompt “focus on the man on the right” to specify the desired focus point. As shown in Fig. 9, Gemini-3 Nano Banana Pro can change the focus to some extent, but it also alters the facial expressions and appearances of people in both the foreground and the background. In contrast, our method preserves the original content while producing a realistic bokeh effect and accurate refocusing.
Figure 10: Visual Comparison for BokehNet Training Ablation.
Appendix F Failure Cases
We observe two main types of unsatisfactory refocusing results: (1) incorrect hallucinated details and (2) inaccurate bokeh blur. For the first type, even though DeblurNet can generate visually plausible high-frequency details, in cases of severe input blur it may hallucinate incorrect detail (see Fig. 11). For the second type, the bokeh blur can be imperfect because our defocus map is derived from monocular depth estimator [7]. When the estimated depth scale is biased, the resulting defocus map may fail to fully cover the intended focus region or may assign too much or too little blur to certain areas. Although our semi-supervised training strategy for BokehNet mitigates these issues to some extent, residual discrepancies with respect to the ground truth can still be observed in challenging cases.
Figure 11: Failure case. Due to the severely blurred input, the model hallucinates an incorrect time, changing 11:30 a.m. to 12:30 a.m.